Hello, Need help finishing my assignment. i am stuck on a domain question?
Find the domain of the following: f(t)= log(t-5) I answered ( is all positive values such that (t-5)>0 or t>5. Is this correct?
g(x)=5e^x I answered all real numbers except 0?
g(x)= ln(x+4) i don't know this one
g(t)=5^t I answered all real numbers except 0
Please help. Thank you.

Answer by jim_thompson5910(21667)

f(t)= log(t-5):
Domain: t>5

--------------

g(x)=5e^x:

Domain: All real numbers (including zero)

You can plug in zero and get a real number as an output (ie it's not undefined).

--------------

g(x)= ln(x+4)

Think of ln(x+4) as log(x+4) and use the same technique as the first problem.

Domain: x>-4

-------------
g(t)=5^t


Domain: All real numbers (including zero)

Once again, you can plug in zero and get a real number as an output.